I have tried to solve this equation, but so far, I have been unsuccessful. I keep getting different answers because I don't know if I am reading the problem correctly. Any help will be greatly appreciated.
The problem given is:
A rectangular field has an unknown width. The length is 2 more yards than the width and the total is area is 120 yards with an exponent of 2 above and to the right the "s" in yards. I am not sure if this means yards squared or square yards.
I have tried either squaring the 120 yds. or finding the square root of 120. In either scenario can I get the numbers to total 120 when plugging the values in for "x".

Answer by rapaljer(4671)

Let x= width of the rectangular field
x+2 = length of the field.

Area = Width * Length
120 = x(x+2)
120 = x^2 +2x

This is a quadratic equation so set it equal to zero:
0=x^2 + 2x - 120

Let's hope it factors:
0=(x-10)(x+12)
x=10 or x=-12

Reject the x=-12, since you can't have a negative width.
x=10 yds = Width
x+2= 12 yds = Length

R^2